# WorldPrompter: Traversable Text-to-Scene Generation

### v1

@@ 3.2 Video Generator for 360° Panoramic Captures
Training a generative model for 360° videos is challenging due to limited data; therefore, we start with an existing video generator. We leverage a pretrained text-to-video diffusion model (similar to [7]) as backbone, and fine-tune it using panoramic data.
@@
Specifically, our model’s backbone architecture $f_{\theta}$ is based on Diffusion Transformers [31]. Typically, in DiT models, a video sequence $x\in\mathbb{R}^{T\times 3\times H\times W}$ is first tokenized into patches, yielding a set of visual tokens $\hat{x}\in\mathbb{R}^{K\times D}$, where $T$, $H$, and $W$ correspond to the temporal and spatial dimensions of the video, $K$ represents the total number of tokens, and $D$ denotes the feature dimension. Being a text-conditioned model, it requires text prompts $c$ that we encode using a pre-trained model [34] and map into the same feature space $D$ by an embedding layer. The transformer $f_{\theta}(\hat{x}_{t};c,t)$ operates on these inputs, denoising the tokens at each timestep $t$. The denoised tokens $\hat{x}_{0}\in\mathbb{R}^{K\times D}$ are subsequently decoded and reassembled into the video data $x_{0}\in\mathbb{R}^{T\times 3\times H\times W}$ via linear layers. To reduce the computational complexity caused by the high-dimensionality of video data, the model performs its denoising in latent space after encoding by a Variational AutoEncoder (VAE) [31, 35].
The scarcity of 360° video data presents a significant challenge in training this model. To develop our high-quality 360° video generator, we curate two carefully selected datasets: (1) a 360° video dataset capturing static indoor environments and (2) a 360° image dataset featuring diverse real-world scenes.
3.2.1 360° video dataset
@@
We collect a 360° video dataset with approximately 1,700 panoramic videos. Each video is captured by a person holding a Ricoh Theta Z1 camera walking inside a nearly static indoor scene to ensure the quality of 3D scene reconstruction. Each video has around 50 seconds of capture, yielding roughly 1,500 frames. To build a dataset for training, we segment the longer video sequences into smaller chunks, typically a 128-frame video clip (5.3/10.6 second at 24/12fps). We use LLaVA [23] to generate detailed captions describing each scene.
3.2.2 360° image dataset
@@
The 360° video data is primarily captured indoors to ensure a static environment with no moving objects during the capture process. However, this may bias the fine-tuned model, limiting its ability to generalize effectively to outdoor environments. To mitigate this issue, we gather another dataset with around 200,000 panoramic images taken with cameras on a tripod, without the photographer in the scene. We caption each image with Blip-2 [19] to describe the whole scene in detail. These stationary images, captured in-the-wild, represent a much broader variety of scenes, thereby enhancing the generalization capability of our video diffusion model.
3.2.3 Mixed Training with Masked Loss
@@
To achieve this, we preprocess all 360° videos by masking out the photographer using a pretrained image segmentation model [1] on each frame in a video. For a 128-frame chunk, we merge all segmentation masks to minimize segmentation errors and ensure a clean region without the person. Additionally, we conservatively mask out the bottom region of the equirectangular panoramic frames to eliminate visible camera equipment and the photographer’s hand. Assuming the corresponding binary mask $M$ is resized into the latent space by nearest neighbor interpolation, the masked diffusion loss is computed as
$\mathbb{E}_{x\sim p_{\text{data}},t\sim U(0,1)}\left[\left\lVert M\odot\left(\epsilon_{t}-f_{\theta}(\hat{x}_{t};c,t)\right)\right\rVert^{2}\right],$ (1)
@@
Our video generator is finally trained on a mix of $1/3$ image data and $2/3$ video data, achieving an empirically reasonable balance. The video diffusion model generates a 128-frame video clip at 12 fps, equivalent to a 10.6-second 360° capture video at a native resolution of 352$\times$704 pixels. We further apply a pretrained GAN-based video upsampler [43] to super-resolve our generated videos 4 times, resulting a final 360° video of 1416x2832 resolution.
3.3 Reconstruction from “Generated” Captures
Given the generated 360° capture video with 128 panoramic frames, we reconstruct a 3D Gaussian scene. For each panoramic frame, we randomly perform 3 perspective crops with a field of view of 120° and resolution of 512$\times$512, resulting in 384 perspective images for pose estimation and reconstruction. Similar to standard scene reconstruction, we calibrate the camera poses of these perspective images using COLMAP [36]. The high average matching ratio $94.6\%$ from COLMAP estimation demonstrates the strong view consistency of the generated videos from our text-to-360° video generator. We adopt Long-LRM, the state-of-the-art feed-forward 3DGS reconstructor [56] to reconstruct the final 3D scene using the estimated camera poses and generated views. Using Long-LRM allows us to achieve fast reconstruction compared to an optimization-based approach. We randomly subsample 32 random views along the original camera path as the inputs since the Long-LRM has a maximum limit for input views. The calibrated camera poses from COLMAP are converted into Plücker rays as the inputs for long-LRM.
3.4 Implementation Details
We fine-tune a pretrained text-to-video diffusion model using mixed 360° image and video datasets. The fine-tuning is conducted on 32 NVIDIA A100 GPUs. 8 GPUs train on 360° images, while 24 generate 128-frame 360° videos. Of these, 12 GPUs produce videos at 24 fps, and 12 generate videos at 12 fps. The model was finetuned from a pretrained text-to-video checkpoint for 20,000 steps in around 200 hours. During inference, we generate videos at 12 fps to extend the traversal range while maintaining good temporal consistency comparable to 24 fps generation.
The inference of a 10.6 second video takes around 10 minutes with an unoptimized PyTorch runtime [30] using 50 diffusion steps, together with video upsampling process. The entire reconstruction takes around 5 minutes, including pose estimation and Long-LRM reconstruction. Our full 3D generation pipeline requires approximately 15 minutes to generate a detailed GS scene, with potential for further optimization to improve inference speed and efficiency for different components.
@@ 3.4 Implementation Details
360DVD [42] 0.7643 3.5797 NA
DreamScene360 [55] 0.8116 4.0942 2.4125
Ours 0.7387 4.2444 2.8043
Table 1: We quantitatively evaluate our method against the state-of-the-art (SOTA) 360° video generation model [42] and the SOTA text-to-scene model [55], which generates 360° coverage scenes. CLIP distance and Q-Align metrics are computed on the generated 360° images or frames across 320 test text prompts. Additionally, Q-Align NVS assesses the quality of novel view synthesis.
A calm Zen rock garden in an ancient monastery courtyard.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
A candlelit medieval cathedral crypt with carved statues.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
A Celtic stone circle on rolling green hills at dawn.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
A glacier-capped polar landscape with drifting icebergs.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
A cramped submarine cafeteria with portholes showing ocean depths.
Refer to caption (a) Panorama Reference Refer to caption (b) Novel View 1 Refer to caption (c) Novel View 2 Refer to caption (d) Novel View 3 Refer to caption (e) Novel View 4
Figure 3: Scene generation results. We present our scene generation results by showing (1) the first 360° image from our generated panoramic video as reference, and (2) novel video renderings (Novel View 1-4) from different spatial locations and viewpoints, demonstrating that our approach effectively synthesizes a traversable, 360° viewable scene. More rendered videos and navigable Gaussian Splats scenes can be found in the supplementary documents.
A cottage in the woods where the walls are made of interwoven tree roots, and the floor is a carpet of moss.
DreamScene360
[55]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours
A cottage surrounded by fields of wildflowers that change color with the seasons, glowing faintly at twilight.
DreamScene360
[55]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours
(a) Panorama Reference
(b) Novel View 1
(c) Novel View 2
(d) Novel View 3
(e) Novel View 4
Figure 4: Comparison of scene generation results. We compare our method with DreamScene360 [55] for text-based scene generation, as both approaches use panoramas as an intermediate representation for generating 360° viewable scenes. The first column presents the panorama reference: for DreamScene360, it is a single generated panorama from a text-to-image model, whereas in our results, it is the first frame of our generated 360° panorama video. While the previous work can produce a 360° viewable scene, its range of motion is highly limited. When novel viewpoints (Novel View 1-4) deviate significantly from the initial position, their GS scene exhibits severe blurriness and artifacts. Besides, our method achieves better alignment with the input prompt and produces the correct global structure of the scene.
4 Results
@@
We demonstrate that our pipeline achieves state-of-the-art quality in both 360° video generation (Sec. 4.1) and 3D scene generation (Sec. 4.2), offering significantly improved visual consistency and view coverage. More visual examples, including videos and Gaussian Splats with an interactive viewer, can be found in the supplementary material. To quantitatively evaluate our method, we prompt ChatGPT [28] to generate a list of 320 descriptions of 3D scenes. We then use CLIP distance [34] to assess prompt adherence and Q-Align [46] to evaluate visual quality.
@@
Additionally, we perform ablation studies in Sec. 4.3 to evaluate our design choices, including mixed data training and masked loss, to ensure the generation of high-quality videos suitable for reconstruction.
4.1 Text-to-360°-Video Generation
@@
Our generation model can generate temporally consistent 360° panoramic videos, exceeding the quality of the state-of-the-art panoramic video generation models. We compare our 360° video generation results with 360DVD [42] in Fig. 5. We showcase four evenly sampled frames comparing 360DVD and our 360° panorama video generation model. Our model generates videos that closely match the input text prompt and produce richer scene details compared to 360DVD. The higher average CLIP distance and Q-Align metric for scenes generated from the 320 prompts in Table 1 further demonstrate the superior prompt alignment and visual quality of our method.
@@
Additionally, videos generated by our model cover a broader scene span compared to 360DVD, which exhibits only slight rotations and translations of the viewpoint as indicated in the marked green box (Fig. 5). More importantly, since the primary goal of 360° panoramic video synthesis is to generate a coherent 3D scene, the quality of 360DVD’s results prevent it from effectively accomplishing this task. Its outputs lack structural consistency across views, exhibiting continuous morphing and noticeable changes, which undermine the perception of a stable 3D environment.
@@
To validate this statement, we run COLMAP on their generated panoramas using a similar cropping and pose estimation procedure as described in Sec. 3.3. We find that the failure rate of COLMAP reaches $89.7\%$, and even in cases where COLMAP succeeds, the match rate remains very low, with only $41.8\%$ images being successfully matched, making their results unsuitable for 3D reconstruction. In contrast, our model preserves view-consistent local structures across frames. We do not encounter any failure cases when running COLMAP, and the average image match rate is as high as $94.6\%$ across our test prompts. This is the key reason we achieve a well-reconstructed 3D scene.
Ours DreamScene360 [55] LucidDreamer [10]
View Coverage 360° 360° Wide-angle front view
Movable Range Large Very Limited Limited
Table 2: Our method achieves state-of-the-art view coverage and traversable range compared to previous 3D generation methods.
4.2 Text-to-Scene Generation
@@
In Fig. 4, we compare our results to the previous state-of-the-art text-to-scene method that can also create 360° virtual environment. However, as an approach that reconstructs a scene from only a single panorama, DreamScene360 cannot create a truly navigable scene. When synthesizing novel views (Novel View 1-4 in Fig. 4) from locations different from the initial position, the rendered images exhibit significant artifacts and blurriness. In Table 1, we compute the Q-align metrics on the rendered novel views (Q-align NVS), where our results achieve higher visual quality and less blurriness compared to previous work, demonstrating that our generated scene enables a more realistic navigation experience. Similarly, we also report the CLIP distance and Q-Align metrics on the generated panorama images, demonstrating that our method generates a 3D environment that more accurately responds to user requests while achieving superior visual quality, even at the initial position.
In addition to quantitative measurements, we highlight that our method is the first to achieve traversable and 360° viewable text-to-scene generation. As shown in Table 2, compared to previous work, our method achieves a complete 360° viewing area while also allowing users to walk inside the generated 3D scene for a large range.
4.3 Ablation Study

### v4

@@ 3.2. Video Generator for 360° Panoramic Captures
Training a generative model for 360° videos from scratch is challenging; therefore, we start with an existing video generator. We leverage a pretrained text-to-video diffusion model as backbone, and fine-tune it using panoramic data.
@@
Specifically, our model’s backbone architecture $f_{\theta}$ is based on Diffusion Transformers (Peebles and Xie, 2023). Typically, in latent DiT models, a video or an image is first encoded by a 3DVAE (Yu et al., 2024b), resulting a latent visual sequence $x\in\mathbb{R}^{T\times 3\times H\times W}$ ($T=1$ if it’s an image). The visual sequence is tokenized into patches, yielding a set of visual tokens $\hat{x}\in\mathbb{R}^{K\times D}$, where $T$, $H$, and $W$ correspond to the temporal and spatial dimensions of the video, $K$ represents the total number of tokens, and $D$ denotes the feature dimension. Being a text-conditioned model, it requires text prompts $c$ that we encode using a pre-trained model (Radford et al., 2021) and map into the same feature space $D$ by an embedding layer. As a decoder-only transformer architecture, the embedded text tokens are concatenated with the noised visual tokens. The transformer $f_{\theta}(\hat{x}_{t};c,t)$ operates on these inputs, denoising the visual tokens at each timestep $t$. The denoised tokens $\hat{x}_{0}\in\mathbb{R}^{K\times D}$ are subsequently decoded and reassembled into the visual data $x_{0}\in\mathbb{R}^{T\times 3\times H\times W}$ via linear layers, and finally decoded as videos or images.
To preserve the generalization ability of the finetuned t2v model across diverse prompts, we curate two carefully selected datasets: (1) a 360° video dataset capturing static indoor environments, and (2) a 360° image dataset featuring a wide range of real-world scenes. We also incorporate a masked diffusion loss, which is key to ensuring artifact-free generation.
3.2.1. 360° video dataset
@@
We start by collecting a 360° video dataset with approximately 1,700 panoramic videos. Each video is captured by a person holding a Ricoh Theta Z1 camera walking inside a nearly static indoor scene to ensure the quality of 3D scene reconstruction. Each video has around 50 seconds of capture, yielding roughly 1,500 frames. To build a dataset for training, we segment the longer video sequences into smaller chunks, typically a 128-frame video clip (5.3/10.6 seconds at 24/12fps). We use LLaVA (Liu et al., 2023a) to generate detailed captions describing each scene.
3.2.2. 360° image dataset
@@
However, as the 360° video data is primarily captured indoors to ensure a static environment with no moving objects during the capture process, this may bias the fine-tuned model, limiting its ability to generalize effectively to outdoor environments. To mitigate this issue, we gather another dataset with around 200,000 panoramic images taken with cameras on a tripod, without the photographer in the scene. We caption each image with Blip-2 (Li et al., 2023a) to describe the whole scene in detail. These stationary images, captured in-the-wild, represent a much broader variety of scenes, thereby enhancing the generalization capability of our video diffusion model.
3.2.3. Mixed Training with Masked Loss
@@
To achieve this, we preprocess all 360° videos by masking out the photographer using a pretrained image segmentation model (Adobe, 2025) on each frame in a video. For a 128-frame chunk, we merge all segmentation masks to minimize segmentation errors and ensure a clean region without the person. Additionally, we conservatively mask out the bottom region of the equirectangular panoramic frames to eliminate visible camera equipment and the photographer’s hand. Assuming the corresponding binary mask $M$ is resized into the latent space by nearest neighbor interpolation, the masked diffusion loss is computed as
(1) $\mathbb{E}_{x\sim p_{\text{data}},t\sim U(0,1)}\left[\left\lVert M\odot\left(\epsilon_{t}-f_{\theta}(\hat{x}_{t};c,t)\right)\right\rVert^{2}\right],$
@@
3.3. Reconstruction from “Generated” Captures
Given the generated 360° capture video with 128 panoramic frames, we reconstruct a 3D Gaussian scene. For each panoramic frame, we randomly perform 3 perspective crops with a field of view of 120° and resolution of $\times512$, resulting in 384 perspective images for pose estimation and reconstruction. Similar to standard scene reconstruction, we calibrate the camera poses of these perspective images using COLMAP (Schönberger and Frahm, 2016). The high average matching ratio $94.6\%$ from COLMAP estimation demonstrates the strong view consistency of the generated videos from our text-to-360° video generator. We adopt Long-LRM, the state-of-the-art feed-forward 3DGS reconstructor (Ziwen et al., 2024) to reconstruct the final 3D scene using the estimated camera poses and generated views. Using Long-LRM allows us to achieve fast reconstruction compared to an optimization-based approach. We randomly subsample 32 random views along the original camera path as the inputs since the Long-LRM has a maximum limit for input views. The calibrated camera poses from COLMAP are converted into Plücker rays as the inputs for long-LRM.
3.4. Implementation Details
We fine-tune a pretrained text-to-video diffusion model using mixed 360° image and video datasets. The fine-tuning is conducted on 32 NVIDIA A100 GPUs with a mixed fps training schedule (12fps and 24fps). The model was finetuned from a pretrained text-to-video checkpoint for 20,000 steps in around 200 hours. During inference, we generate videos at 12 fps to extend the traversal range while maintaining good temporal consistency comparable to 24 fps generation.
The inference of a 10.6-second video takes around 10 minutes with an unoptimized PyTorch runtime (Paszke et al., 2019) using 50 diffusion steps, together with the video upsampling process. The entire reconstruction takes around 5 minutes, including pose estimation and Long-LRM reconstruction. Our full 3D generation pipeline requires approximately 15 minutes to generate a detailed GS scene, with potential for further optimization to improve inference speed and efficiency for different components.
Figure 3. Comparison of 360° video generation results. We present a comparison of 360° panoramic video generation results against 360DVD (Wang et al., 2024). Our generated videos exhibit significantly better visual quality and prompt alignment. We manually mark an anchor region with a green box in the 360° image, highlighting the substantially longer traversal range of our generated videos. Stronger parallax is important for accurate 3D reconstruction.
Figure 4. We conduct an ablation study to evaluate the impact of the masked diffusion loss. This loss function effectively guides the model to generate only the pixels of the scene, eliminating the moving photographer present in the original video training data. As a result, it enables the generation of a complete 360° scene capture for reconstruction.
4. Results
@@
By generating 360° video capture, we demonstrate that our pipeline achieves state-of-the-art quality in 3D scene generation (Sec. 4.2), offering significantly improved view coverage, traversability and coherent global visual content. This is primarily due to a strong 360° video generator for scenes (Sec. 4.1). To quantitatively evaluate our method, we prompt ChatGPT (OpenAI, 2023) to generate a list of 320 descriptions of 3D scenes. Using this in-the-wild test set, we conduct systematic comparisons against existing methods, demonstrating superior video and 3D scene generation quality through both visual examples and quantitative metrics. More visual examples can be found in the supplementary material.
@@
Additionally, we perform ablation studies in Sec. 4.3 to evaluate our design choices, including mixed data training and masked loss, to ensure the generation of high-quality videos suitable for artifact-free 3D scene reconstruction.
Table 1. We quantitatively evaluate our method against the state-of-the-art (SOTA) 360° video generation model (Wang et al., 2024). We report CLIP distance (CLIP) for prompt alignment quality, Q-Align score for image quality, VBench for aesthetic/imaging quality, and COLMAP Failure Rate (COLMAP-FR) along with Matching Rate (COLMAP-MR) for multi-view consistency. The metrics are computed on the generated 360° frames across 320 test text prompts. Quantitative metrics across multiple dimensions demonstrate that our generated 360° videos achieve high visual quality and can be reliably used for 3D scene representation and reconstruction.
[…]
360DVD 0.7643 0.6704 0.538/0.661 89.7% 41.8%
Ours 0.7387 0.8334 0.615/0.717 0% 94.6%
Table 2. We quantitatively evaluate our method against several text-to-scene models, including (Chung et al., 2023), (Zhou et al., 2024), and (Yang et al., 2024). To assess prompt adherence and visual quality, we compute CLIP distance and Q-Align metrics on the generated 360° images or frames across 320 test prompts randomly generated by ChatGPT. Note that LucidDreamer is an image-to-scene generation model which requires both text and image as inputs, and we generate the initial image using ChatGPT; therefore CLIP and Q-Align score evaluating the panorama synthesis quality is not applicable for LucidDreamer. Our method outperforms all baselines across these evaluations. Furthermore, we conduct a 4AFC user study to assess three key aspects of scene quality: Realism/Completeness/Traversal Consistency and report the selection percentages.
[…]
LucidDreamer N/A N/A 21.0% 10.0% 10.0%
DreamScene360 0.8116 0.8074 7.0% 6.0% 6.0%
LayerPano3D 0.7977 0.8312 34.0% 25.0% 29.0%
Ours 0.7387 0.8334 38.0% 59.0% 55.0%
4.1. Text-to-360°-Video Generation
@@
A 360° video generator that accurately reproduces the distribution of real-world video captures is key to making it a sufficient proxy for 3D scene representation. Our generation model can generate spatial-temporal consistent 360° panoramic videos, exceeding the quality of the state-of-the-art panoramic video generation models. We compare our 360° video generation results with 360DVD (Wang et al., 2024) in Fig. 3. We showcase four evenly sampled frames comparing 360DVD and our 360° panorama video generation model. Our model generates videos that closely match the input text prompt and produce more realistic scene details compared to 360DVD. In Table 1, we evaluate generation quality across multiple axes, where our generator consistently outperforms previous SOTA in common image/video generation quality metrics e.g., CLIP (Radford et al., 2021), Q-align (Wu et al., 2023), VBench (Huang et al., 2024).
@@
Additionally, videos generated by our model cover a broader scene span compared to 360DVD, which exhibits only slight rotations and translations of the viewpoint as indicated in the marked green box (Fig. 3). More importantly, since the primary goal of 360° panoramic video synthesis is to generate a coherent 3D scene, the quality of 360DVD’s results prevents it from effectively accomplishing this task. Its outputs lack structural consistency across views, exhibiting continuous morphing and noticeable changes, which undermine the perception of a stable 3D environment.
@@
To validate this statement, we run COLMAP on their generated panoramas using a similar cropping and pose estimation procedure as described in Sec. 3.3. As shown in Table 1, we find that the failure rate of COLMAP reaches $89.7\%$ (COLMAP-FR), and even in cases where COLMAP succeeds, the match rate (COLMAP-MR) remains very low, with only $41.8\%$ images being successfully matched, making their results unsuitable for 3D reconstruction. In contrast, our model preserves view-consistent local structures across frames. We do not encounter any failure cases when running COLMAP, and the average image match rate is as high as $94.6\%$ across our test prompts. This is the key reason our 360° videos can be a reasonable 3D representation for reconstruction.
4.2. Text-to-Scene Generation
@@
In Fig. 7, we compare our results to the previous state-of-the-art text-to-scene method that can also create 3D scenes. LucidDreamer (Chung et al., 2023), an outpainting-based scene generation method, synthesizes single-perspective views using an image diffusion model, but suffers from global geometric distortions and lacks support for full-view navigational experiences. Other methods that construct 3D scenes from panoramic images—such as DreamScene360 (Zhou et al., 2024) and LayerPano3D (Yang et al., 2024)—provide 360° view coverage, but do not enable truly navigable scenes, as significant camera movement often breaks scene coherence. For all other methods, when synthesizing novel views (Novel View 1–4 in Fig. 7) from positions different from the initial viewpoint, the rendered images exhibit significant artifacts, including blurriness, missing regions, and structural distortions. Similarly, we report the CLIP distance and Q-Align metrics on the generated 360° images, demonstrating that our method generates a 3D environment that more accurately responds to user requests while achieving superior visual quality, even at the initial position.
Figure 5. We verify that incorporating additional panorama image data helps the model generalize to diverse environments. Without the image data, the generated 360° videos not only exhibit degraded visual appearance but also display messy and incorrect content at the bottom of the video. This issue arises because these pixels lack supervision due to the masked loss.
We argue that quality assessment is best conducted within the context of a real navigation experience. To this end, we perform a user study in which participants are presented with navigation videos generated by different methods. For each approach, we render a short video clip simulating camera movement through the scene, and ask users to select the most preferred result. Participants are instructed to evaluate the methods using a four-alternative forced-choice (4AFC) paradigm for the following dimensions: (1) Realism: Realistic Visual (2) Completeness: No missing regions (3) Traversal Consistency: Stable, distortion-free geometry during navigation. We collect 20 responses and report the selection percentages in Table 2. Our method received the majority of votes across all evaluation criteria, with particularly strong margins in Completeness and Traversal Consistency.
4.3. Ablation Study